OpenStudy (anonymous):

* Logx5=4log5(X) find solution

OpenStudy (ghazi):

use base change formula

OpenStudy (anonymous):

\[5^{\pm \frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

Kindly solve it...

OpenStudy (ghazi):

log5/logx= 4(log x/log5)

hartnn (hartnn):

is it \[\log_{}x^5=4\log_{}5x \] ?

OpenStudy (ghazi):

\[\log^25=4*\log^2x....2\log5=8logx .......x=1/4\log5\]

OpenStudy (ghazi):

\[logx= \frac{ 1 }{ 4 }\log 5\]

OpenStudy (ghazi):

\[x=5^{1/4}\]

hartnn (hartnn):

how did u get log x=1/4 log5 from x=1/4 log 5 ??

OpenStudy (ghazi):

well that was a typing mistake my apology :)

OpenStudy (ghazi):

after that i haven't repeated it

hartnn (hartnn):

then x won't be 5^(1/4)

OpenStudy (ghazi):

yes it will be

hartnn (hartnn):

but x=1/4 log 5,how can it be 5^1/4

hartnn (hartnn):

it can be atmost log(5^(1/4))

OpenStudy (ghazi):

\[\log x= \frac{ 1 }{ 4 } \log5......\log x =\log 5^{1/4}....x=5^{1/4}\]

hartnn (hartnn):

ok.

OpenStudy (ghazi):

:)